# Improving the Euclidean Diffusion Generation of Manifold Data by Mitigating Score Function Singularity

### v1

@@ 7 Related work
It is important to clarify that research on manifold-related diffusion models can be broadly categorized into two directions: (1) diffusion models for distributions under the manifold hypothesis, where the underlying manifold is unknown, and (2) diffusion models for distributions on a known manifold, which is the focus of this work. For the former, we discuss the singularity of score functions in related work; for the latter, we explore several studies based on known manifolds.
On the singularity of score functions
@@
Several studies on diffusion models under the manifold hypothesis [23, 21, 6] have highlighted the singularity of the score function, specifically identifying that as $t\to 0$, the score function diverges. Beyond empirical observations, recent studies have provided mathematical analysis of diffusion models based on the VPSDE. For instance, Chen et al. [2] presents a mathematical analysis under a linear assumption on the dataset, where each data point is assumed to lie on a hyperplane. In [23], it is shown that the score function norm satisfies $\mathbb{E}\|\nabla_{x}\log p_{t}(x))\|\gtrsim 1/\sqrt{t}$. Furthermore, Lu et al. [21] rigorously examines the asymptotic behavior of the score function as $t\to 0$ for the VPSDE, demonstrating that its singularity follows a $1/t$ scaling in a strong pointwise sense, which is similar to our results in (10). Further discussion can be found in [19].
@@
To mitigate this singularity issue, a number of works (e.g. [31]) propose starting their analysis at $t=\epsilon>0$, instead of $t=0$. However, this approach is equivalent to selecting a larger $\sigma_{\min}$ and does not fully resolve the fundamental singularity problem. The study [8] introduces a diffusion model on the product space of position and velocity, employing hybrid score matching to circumvent the singularity. Furthermore, in [30], the score function is parameterized by $s_{\theta}(x,t)=\hat{s}_{\theta}(x)/\sigma_{t}$, where $\hat{s}_{\theta}(x)$ is a neural network.
@@
Riemannian Score-based Generative Models [7] extend SDE-based diffusion models to manifold settings by estimating the transition kernels using either the eigenpairs of the Laplacian–Beltrami operator on manifolds [20] or the Varadhan approximation via the Riemannian logarithmic map. Riemannian Diffusion Models [13] employ a variational diffusion framework for Riemannian manifolds and, similar to our approach, consider submanifolds embedded in Euclidean space. Additionally, Trivialized Diffusion Models [37] introduce a technique called trivialization that adapts diffusion models from Euclidean spaces to Lie groups.
@@ 8 Conclusion
We address the multiscale singularity of the score function for manifold-constrained data, which limits the accuracy of Euclidean diffusion models in this work. By decomposing the score function into tangential and normal components, we identify the source of scale discrepancies and propose two methods: Niso-DM, which reduces discrepancies with non-isotropic noise, and Tango-DM, which trains only the tangential component. Both methods achieve superior performance on complex manifolds, improving the accuracy and robustness of diffusion-based generative models.
Appendix A Proof for Theorem 3.1
@@
For simplicity of notation, we use ||$\cdot$|| ⋅ | to denote the Euclidean norm in the following two sections. Let us recall the expression
@@
Let $x^{*}\in\mathcal{M}$ be the unique minimizer of the minimization
$\min_{x\in\mathcal{M}}|x-\tilde{x}|.$ (24)
In the following, we study the limit of integrals over $\mathcal{M}$ as $\sigma\rightarrow 0$ by a change of variables using the exponential map on $\mathcal{M}$ at $x^{*}$. To this end, let us first derive some asymptotic expressions related to the exponential map.
Denote by $T_{x^{*}}\mathcal{M}=\{v|\nabla\xi(x^{*})^{T}v=0,v\in\mathbb{R}^{n}\}$ the tangent space of $\mathcal{M}$ at $x^{*}$. Given $v\in T_{x^{*}}\mathcal{M}$, let $\gamma_{v}(t)\in\mathbb{R}^{n}$ be the geodesic curve on $\mathcal{M}$ starting from $\gamma_{v}(0)=x^{*}$ at $t=0$ such that $\dot{\gamma}_{v}(0)=\mathrm{d}\gamma_{v}(t)/\mathrm{d}t(0)=v$. We have Taylor’s expansion
@@
Since $x^{*}=\gamma_{v}(0)$ is the minimizer of (24), we know that $|\gamma_{v}(t)-\tilde{x}|^{2}$ (as a function of $t$) attains its minimum at $t=0$. Therefore, taking derivative and setting $t=0$, we obtain
@@
An expression for the second-order derivative $\ddot{\gamma}_{v}(0)$ in (25) is not available in general. However, by differentiating the identity $\big{(}I-P(\gamma_{v}(t))\big{)}\dot{\gamma}_{v}(t)=0$ with respect to $t$, where and setting $t=0$, we obtain the equation (i.e. the expression for the orthogonal component of $\ddot{\gamma}_{v}(0)$)
$\big{(}I-P(x^{*})\big{)}\ddot{\gamma}_{v}(0)=\sum_{1\leq j,j^{\prime}\leq n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(x^{*})v_{j}v_{j^{\prime}}\,,$ (27)
@@
$|v|^{2}\ddot{\gamma}_{\hat{v}}(0)\cdot(x^{*}-\tilde{x})=|v|^{2}\Big{(}(I-P(x^{*}))\ddot{\gamma}_{\hat{v}}(0)\Big{)}\cdot(x^{*}-\tilde{x})=\sum_{1\leq i,j,j^{\prime}\leq n}\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}(x^{*})v_{j}v_{j^{\prime}}(x^{*}_{i}-\tilde{x}_{i})\,.$ (30)
@@
$S_{jj^{\prime}}=\delta_{jj^{\prime}}+\frac{1}{2}\sum_{i=1}^{n}\Big{(}\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}\Big{)}(x^{*})(x^{*}_{i}-\tilde{x}_{i})\,,\quad\mbox{for}~{}1\leq j,j^{\prime}\leq n\,.$ (32)
From the Gershgorin circle theorem, all eigenvalues of $S$ are greater than $1-\frac{1}{2}\max_{j}\sum_{i,j^{\prime}}|(\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}})(x^{*})(x^{*}_{i}-\tilde{x}_{i})|$. Furthermore, this lower bound is positive, as $\|x^{*}-\tilde{x}\|_{\infty}<\frac{1}{n^{2}M_{1}}$. Therefore, the matrix $S$ is positive definite. With the above expansions, we prove the two claims.
@@
We consider the general case where $\tilde{x}\in\mathbb{R}^{n}$ may not belong to $\mathcal{M}$. Since $p_{\sigma}$ is the probability density of Gaussian distribution, we can derive
@@
On the one hand, since $x^{*}$ is the unique minimizer of (24), there exists $\delta>0$, such that $|x-\tilde{x}|\geq|x^{*}-\tilde{x}|+\delta$ for all $x\in\mathcal{M}\setminus\mathcal{O}^{\prime}$, where $\mathcal{O}^{\prime}:=\exp_{x^{*}}(\mathcal{O})$. Therefore, we have
@@
$\nabla_{\tilde{x}}p_{\sigma}(\tilde{x})=$ $-\int_{\mathcal{M}}p_{0}(x)\nabla_{x}p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ (41)
@@
where $\nabla^{\mathcal{M}}_{x}$ denotes the gradient operator on $\mathcal{M}$ at $x$ and we used the fact that $\nabla^{\mathcal{M}}_{x}=P(x)\nabla_{x}$ to obtain the third equality, and the last equality follows by using the integration by parts formula on $\mathcal{M}$.
For the first term in the last line of (41), similar to (39), we can obtain

### v3

@@ 7 Related work
Research on manifold-related diffusion models can be broadly categorized into two directions: (1) diffusion models for distributions under the manifold hypothesis, where the underlying manifold is unknown, and (2) diffusion models for distributions on a known manifold, which is the focus of this work. For the former, we discuss the singularity and the scale discrepancy of score functions mentioned in the related work; for the latter, we explore several studies based on known manifolds.
Singularity of score functions.
@@
Several studies on diffusion models under the manifold hypothesis [29, 27, 3] have highlighted the divergence of the score function as $t\to 0$. Beyond empirical observations, recent studies have provided mathematical analysis of diffusion models based on the VPSDE. For instance, Chen et al. [4] present a mathematical analysis under the assumption that each data point lies on a hyperplane. In [29], it is shown that the norm of the score function satisfies $\mathbb{E}\|\nabla_{x}\log p_{t}(x))\|\gtrsim 1/\sqrt{t}$. Furthermore, Lu et al. [27] rigorously prove that this singularity follows a $1/t$ scaling in a strong pointwise sense under the VPSDE, which is similar to our results in (10). Further discussion can be found in [25].
@@
To mitigate this singularity issue, a number of works (e.g. [38]) assume that the initial time is at $t=\epsilon>0$, instead of $t=0$. The study [9] introduces a diffusion model on the product space of position and velocity, employing hybrid score matching to circumvent the singularity. Furthermore, in [37], the score function is parameterized by $s_{\theta}(x,t)=\hat{s}_{\theta}(x)/\sigma_{t}$, where $\hat{s}_{\theta}(x)$ is a neural network.
Scale discrepancy of the score function.
The work [39] uses the singular value decomposition of the score matrix to identify the intrinsic dimension of the manifold, which essentially utilizes the discrepancy of the score function. References [42, 20, 44, 46] investigate the fundamental principles of diffusion models under the manifold hypothesis, by studying the eigen-decomposition of the Jacobian of the score functions. This analysis implicitly relates to the score discrepancy discussed in our work. In particular, the decomposition of the denoiser mapping (Eq. (12)) in [20], which involves a geometry-adaptive harmonic basis, also implies this discrepancy, aligning with (10) in our paper. Papers [1] and [10] study the Memorization and Generalization in Generative Diffusion under the manifold hypothesis through Hidden Manifold Models and Generalized Linear Models. Besides, [16] suggests that a conservative diffusion is guaranteed to yield the correct conclusions when analyzing local features of the data manifold.
While previous works have implicitly touched upon the scale differences in manifold settings, our work explicitly formalizes this phenomenon with rigorous mathematical formulations and introduces novel solutions to address it. Furthermore, we believe our methods can enhance the accuracy of diffusion models during the manifold consolidation phase (as described in [42]) or the collapse transition times (as mentioned in [1, 10]), by refining the computation of the tangential component of the score function.
Additional noise along the normal direction.
Paper [15] proposes adding noise along the normal direction to "inflate" manifolds under the setting of normalizing flows, which is similar to our Niso method. Theorem 4 and Proposition 7 in [15] analyze the relationship between the perturbed distribution and the original distribution on the manifold, which aligns with the conclusions in our equations (12) and (18).
On the other hand, our method differs from [15] in several key aspects. In [15], the added noise has a constant magnitude, requiring additional assumptions (e.g., Q-normally reachable) to prevent interference between noise at different points. In contrast, our analysis (Theorem 3.1) permits the noise magnitude to diminish toward zero, relying on weaker assumptions. Additionally, [15] applies noise inflation prior to training the normalizing flow, whereas our method integrates noise addition directly into the diffusion process, accompanied by the noise dynamics of diffusion models. Finally, while the analysis in [15] focuses on the perturbed density function, our work studies the asymptotic behavior of the score function.
@@
Riemannian Score-based Generative Models [8] extend SDE-based diffusion models to manifold settings by estimating the transition kernels using either the Riemannian logarithmic map or the eigenpairs of the Laplacian–Beltrami operator on manifolds [26]. Riemannian Diffusion Models [17] employ a variational diffusion framework for Riemannian manifolds and, similar to our approach, consider submanifolds embedded in Euclidean space. Additionally, Trivialized Diffusion Models [49] adapt diffusion models from Euclidean spaces to Lie groups.
8 Limitations and future work
Compared to approaches under the manifold hypothesis, our method still requires knowledge of the manifold’s definition. In fields like image and language processing, the manifold structure is often assumed but not explicitly characterized. Future work will focus on extending our approach to scenarios where the manifold is undefined or implicitly represented. Furthermore, Tango-DM is not suitable for the Reverse SDE Sampling algorithm and tends to be slower due to the annealing sampler. Designing Reverse SDE-style algorithms for Tango-DM remains an open direction for future research.
@@ 9 Conclusion
We address the multiscale singularity of the score function for manifold-structured data, which limits the accuracy of Euclidean diffusion models. By decomposing the score function into tangential and normal components, we identify the source of scale discrepancies and propose two methods: Niso-DM, which reduces discrepancies with non-isotropic noise, and Tango-DM, which trains only the tangential component. Both methods achieve superior performance on complex manifolds, improving the accuracy and robustness of diffusion-based generative models.
Acknowledgments and Disclosure of Funding
Zichen Liu acknowledges support from the China Scholarship Council (Grant No. 202306010047). Wei Zhang acknowledges support from DFG’s Eigene Stelle (Project No. 524086759). Tiejun Li acknowledges support from the National Key R&D Program of China (Grant No. 2021YFA1003301) and the National Science Foundation of China (Grant No. 12288101).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of the work are discussed in the Appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the assumptions in the statement of the theorem and the proof of the theorem in the Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The code and the datasets are provided.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide open access to the data and the code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All the training and test details are provided in the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Error bars are provided in the table.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The information on the computer resources is provided in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The potential positive societal impacts can be found in the Appendix and there is no negative societal impact of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators or original owners of all assets used in the paper are properly credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Proof of Theorem 3.1
@@ 80. Guidelines:
For simplicity of notation, we use $|\cdot|$ to denote the Euclidean norm in the following two sections. We assume that $\tilde{x}\in\mathbb{R}^{n}$ is fixed. Let us recall the expression
@@
Let $x^{*}\in\mathcal{M}$ be the unique minimizer of the minimization problem
$x~|.\min_{x\in\mathcal{M}}$|x-$\tilde${x}|. (24)
In the following, we study the limit of the integral in (23) over $\mathcal{M}$ as $\sigma\rightarrow 0$ by a change of variables using the exponential map on $\mathcal{M}$ at $x^{*}$. To this end, let us first derive some asymptotic expressions related to the exponential map.
Denote by $T_{x^{*}}\mathcal{M}=\{v|\nabla\xi(x^{*})^{T}v=0,v\in\mathbb{R}^{n}\}$ the tangent space of $\mathcal{M}$ at $x^{*}$. Given $v\in T_{x^{*}}\mathcal{M}$, let $\gamma_{v}(t)\in\mathbb{R}^{n}$ be the geodesic curve on $\mathcal{M}$ starting from $\gamma_{v}(0)=x^{*}$ at $t=0$ such that $\dot{\gamma}_{v}(0)=v$. We have Taylor’s expansion
@@
Since $x^{*}=\gamma_{v}(0)$ is the minimizer of (24), we know that $|\gamma_{v}(t)-\tilde{x}|^{2}$ (as a function of $t$) attains its minimum at $t=0$. Therefore, taking the derivative and setting $t=0$, we obtain
@@
An expression for the second-order derivative $\ddot{\gamma}_{v}(0)$ in (25) is not available in general. However, by differentiating the identity $\big(I-P(\gamma_{v}(t))\big)\dot{\gamma}_{v}(t)=0$ with respect to $t$, and setting $t=0$, we obtain the equation (i.e. the expression for the orthogonal component of $\ddot{\gamma}_{v}(0)$)
$\big(I-P(x^{*})\big)\ddot{\gamma}_{v}(0)=\sum_{j,j^{\prime}}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(x^{*})v_{j}v_{j^{\prime}}\,$, (27)
@@
$|v|^{2}\ddot{\gamma}_{\hat{v}}(0)\cdot(x^{*}-\tilde{x})=|v|^{2}\Big((I-P(x^{*}))\ddot{\gamma}_{\hat{v}}(0)\Big)\cdot(x^{*}-\tilde{x})=\sum_{i,j,j^{\prime}}\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}(x^{*})v_{j}v_{j^{\prime}}(x^{*}_{i}-\tilde{x}_{i})\,.$ (30)
@@
$S_{jj^{\prime}}=\delta_{jj^{\prime}}+\frac{1}{2}\sum_{i}\Big(\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}\Big)(x^{*})(x^{*}_{i}-\tilde{x}_{i})\,,\quad\mbox{for}~1\leq j,j^{\prime}\leq n\,.$ (32)
From the Gershgorin circle theorem [41], all eigenvalues of $S$ are greater than $1-\frac{1}{2}\max_{j}\sum_{i,j^{\prime}}|(\frac{\partial P_{ij^{\prime}}}{\partial x_{j}}+\frac{\partial P_{ij}}{\partial x_{j^{\prime}}})(x^{*})(x^{*}_{i}-\tilde{x}_{i})|$. Furthermore, this lower bound is positive, as $\|x^{*}-\tilde{x}\|_{\infty}<\frac{1}{n^{2}M_{1}}$. Therefore, the matrix $S$ is positive definite. With the above expansions, we prove the two claims.
@@
We consider the general case where $\tilde{x}\in\mathbb{R}^{n}$ may not belong to $\mathcal{M}$. Since $p_{\sigma}(\tilde{x}|x)$ is the probability density of Gaussian distribution $\mathcal{N}(x,\sigma^{2}I)$, we can derive
@@
On the one hand, since $x^{*}$ is the unique minimizer of (24), there exists $\delta>0$, such that $|x-\tilde{x}|^{2}\geq|x^{*}-\tilde{x}|^{2}+\delta$ for all $x\in\mathcal{M}\setminus\mathcal{O}^{\prime}$, where $\mathcal{O}^{\prime}:=\exp_{x^{*}}(\mathcal{O})$. Therefore, we have
@@
[…]
@@
where $\nabla^{\mathcal{M}}_{x}$ denotes the gradient operator on $\mathcal{M}$ at $x$ and we used the fact that $\nabla^{\mathcal{M}}_{x}=P(x)\nabla_{x}$ to obtain the third equality, and the last equality follows by using the integration by parts formula on $\mathcal{M}$. For the first term in the last line of (41), similar to (39), we can obtain